Troubleshooting: fluorescence module

From Bioblast
Jump to: navigation, search

Header ecosystem 2020.jpg

OROBOROS support
        MitoPedia: O2k hardware         MitoPedia: DatLab         MitoPedia: SUIT         O2k-Procedures         O2k-Manual         O2k-Videosupport         MitoPedia: O2k-Open Support        


Troubleshooting: fluorescence module


Template NextGen-O2k.jpg
O2k-Open Support


MitoPedia O2k and high-resolution respirometry: O2k-Open Support 



In the following page, the most common problems occurring with either the O2k-Fluo Smart-Module or O2k-Fluo LED2-Module are summarized. The problem can be related to (a) the fluorescence settings in DatLab, (b) the fluorescence sensors, (c) the Fluorescence-Control Unit (in the case of the O2k-Fluo LED2-Module), or (d) the O2k.

Zero fluorescence signal

Problem: The raw fluorescence signal is 0 V with and without the fluorescence dye.
Solution: To identify cause of the problem, please go through the following check-list:
1. Check that the Fluo-Sensors are connected to the O2k and inserted into the O2k-chamber in the right position. See: O2k-Fluo_Smart-Module#Connect_Fluorescence-Sensor_to_O2k.
2. Check that the Fluorescence-Control Unit is properly connected to the O2k of series D-G.
3. In DatLab check that:
a) the illumination is switched off in the ´O2k control´window.
b) the gain for Fluo-Sensor and the Fluo intensity are set to a number greater than 0 in the 'Amperometric,Amp' tab of the ´O2k control´ window.
4. If all the above-mentioned settings and set up are correct and the detected signal with the Fluo-Sensor is still 0, perform the following test to localize the problem:
Add distilled water or respiration medium into the O2k-chamber (without sample), insert the black stopper and change the settings for fluorometric measurements: Gain=1000; Fluo intensity (Serie H or higher) or Polarisation voltage (Series D-G)=500.
  • 1. run: Fluo-Sensor A in chamber A and Fluo-Sensor B in chamber B (let it run for 5 min)
  • 2. run: Fluo-Sensor A in chamber B and Fluo-Sensor B in chamber A (let it run for 5 min)
If the problem is with the Fluo-Sensor, the fluorescence signal will be 0 V with the faulty Fluo-Sensor in both chambers. In this case, the Fluo-Sensor should be shipped to us for repairs.
If the problem is with the O2k or the O2k-LED2 Fluo-Module, 0 V fluorescence signal can be observed just in one chamber with two different Fluo-Sensors. In this case, the O2k or the Fluorescence-Control Unit (Series D-G) should be shipped to us for repairs. In some cases, the defect of the Fluorescence-Control Unit (Series D-G) leads to 0 V in both chambers.
Please contact our O2k-technical Support to analyze your problem before shipment. Be sure to always send us original DatLab files with the above-mentioned tests to evaluate Fluo-Sensor performance.

Noisy fluorescence signal

Problem: The raw fluorescence signal is noisy to such an extent that the evaluation of the results is not possible.
Solution: To localize the problem, perform the following test:
  • Add distilled water or respiration medium into the O2k-chamber (without sample), insert the black stopper and change the settings for fluorometric measurements: Gain=1000; Fluo intensity (Serie H or higher) or Polarisation voltage (Series D-G)=500.
  • 1. run: Fluo-Sensor A in chamber A and Fluo-Sensor B in chamber B (let it run for 5 min)
  • 2. run: Fluo-Sensor A in chamber B and Fluo-Sensor B in chamber A (let it run for 5 min)
If the problem is with the Fluo-Sensor, the fluorescence signal will be noisy with the faulty Fluo-Sensor in both chambers. In this case, the Fluo-Sensor should be shipped to us for repairs.
If the problem is with O2k or the O2k-LED2 Fluo-Module, the fluorescence signal is noisy in one chamber with two different Fluo-Sensors (see example below). In this case, the O2k or the Fluorescence-Control Unit (Series D-G) should be shipped to us for repairs. If the noise occurs in both chambers, the Fluo-Sensors and the O2k (and Fluorescence-Control Unit - series D-G) should be shipped to us.
Please contact our O2k-technical Support to analyze your problem before shipment. Be sure to always send us original DatLab files with the above-mentioned tests to evaluate Fluo-Sensor performance.
Example: Problem with O2k or O2k-LED2 Fluo-Module
Provided by Hazirah Hassan, University Kebangsaan Malaysia, hazy_hassan@yahoo.com
Hassan Hazirah 1.test.png
Hassan Hazirah 2.test.png
Hassan Hazirah biological sample.png
After changing the Fluo-Sensors between the O2k-chambers, the problem is related to the O2k or O2k-LED2 Fluo-Module since the fluorescence signal is negative and unstable in chamber B with both Fluo-Sensors. In the third figure, the fluorescence signal is very unstable in chamber B, which makes the experiment hard to interpret. Both the O2k and the LED2 Fluo-Module should be shipped to our electronic workshop.

Negative fluorescence signal

Problem: The fluorescence signal of one Fluo-Sensor is negative.
Solution: To localize the problem, perform the following test:
  • Add distilled water or respiration medium (without sample) into the O2k-chamber, insert the black stopper, and change the settings for fluorometric measurements: Gain=1000; Fluo intensity (Series H or higher) or Polarisation voltage (Series D-G) = 500.
  • 1. run: Fluo-Sensor A in chamber A and Fluo-Sensor B in chamber B (let it run for 5 min)
  • 2. run: Fluo-Sensor A in chamber B and Fluo-Sensor B in chamber A (let it run for 5 min)
The problem is related to one of the Fluo-Sensors, if you observe a negative signal in both chambers after swapping them. The negative signal might be related to a wet plug of the Fluo-Sensor. The solution is to dry it out with e.g. a hair dryer and let it run connected to the O2k (25 °C, illumination off, distilled water in the O2k-chamber, gain for Fluo-Sensor 1000, fluorescence intensity=500) for a couple of days in a dry and warm environment. If this treatment does not solve the problem and you get still negative fluorescence values, the Fluo-Sensor should be sent back to us for further investigation.
Please contact our O2k-technical support to analyze your problem before shipment. Be sure to always send us original DatLab files with the above-mentioned tests to evaluate Fluo-Sensor performance.
Example: Provided by Anne Laure Charles, FR_Strasbourg_Zoll J,anne laure charles <anne.laure.charles@outlook.fr
Anne Laure Charles files.png
The fluorescence signal was negative in both chambers after swapping the Fluo-Sensors, but the signal changed after swapping the sensors, which means that the problem is located to the Fluo-Sensors. In this case, both Fluo-Sensors should be shipped to the electronic workshop.

Fluorescence signal is constantly at 10 V

Problem: The fluorescence signal is constantly at the 10 V upper limit when using the O2k-Fluo Smart-Module or O2k-Fluo LED2-Module.
Solution: This is probably due to high light exposure to the photodiode. Check the following parameters:
1. The chamber illumination must be switched off. Check it in the O2k control window.
2. The photodiode filters must be inserted, otherwise the light emitted from the LED will not be filtered. Information on how to mount the filter cap- » here.
3. Black PEEK stoppers must be used.
4. The fluorescence intensity cannot be set too high. Check in the Amperometric,Amp tab of the ´O2k control´ window.
If all the above-mentioned settings and set up are correct and the detected signal is still at 10 V (out of scale), the O2k (and Fluorescence-Control Unit - series D-G) or/and the Fluo-Sensors should be shipped to us.
Please contact our O2k-technical Support to analyze your problem before shipment. Be sure to always send us original DatLab files with the above-mentioned tests to evaluate Fluo-Sensor performance.
Questions.jpg

Useful links

» Amplex UltraRed
» Safranin
» TMRM
» Rhodamine 123
» Magnesium Green
» Calcium Green
» O2k-FluoRespirometer manual
» O2k-Fluo LED2-Module
» O2k-Demo experiments
» Sole source information
» Oroboros O2k